OpenStudy (anonymous):

Having problems solving this rational equation. Can anyone help? 30 50 1 ---- = ----- + ---- X x+10 2

OpenStudy (helder_edwin):

multiply the equation by 2x(x+10) tell me what you get

OpenStudy (anonymous):

2x(x+10) is the LCD?

OpenStudy (helder_edwin):

yes

OpenStudy (anonymous):

I get lost in the steps. I have no problems solving this question below. Is there anyway you could show me the original equation? 1 1 --- = ---- x 6-x

OpenStudy (anonymous):

What I got was -30 for the first question, which I know is wrong.

OpenStudy (helder_edwin):

OK let's see. your original equatio was (it is at the top) \[ \large \frac{30}{x}=\frac{50}{x+10}+\frac{1}{2} \] multiplying by 2x(x+10) we get \[ \large 2\cdot30(x+10)=2\cdot50x+x(x+10) \] \[ \large 60x+600=100x+x^2+10x \] \[ \large 0=x^2+50x-600 \] \[ \large 0=(x+60)(x-10) \]

OpenStudy (anonymous):

Thank you. I got the left side down to 60x + 600. I didn't understand to move both and make it 0. Thank you

OpenStudy (helder_edwin):

u r welcome